A wholesaler of pencil charges 24/- per dozen on an order of fifty dozen or less. For order in excess of fifty dozen, the price is reduced by 20 paise per dozen. Find the size of the order in order to maximize revenue
1
Expert's answer
2018-09-28T10:52:08-0400
When x>50, P = 24 – 0,2*(x-50) = 34 – 0,2x
Thus, TR = P*X = (34-0,2*X) = 34*x – 0,2*x^2
For maximum TR, we have d(TR)/dx = 34- 0,4*x = 0, or x=34/0,4 = 85 dozens.